The length of a rectangle plus its width is 18 cm. The area is 80 square cm. what are the length and width of the rectangle?
Answer by checkley77(12844)

L+W=18
L=18-W
LW=80
(18-W)W=80
18W-W^2=80
W^2-18W+80=0
(W-10)(W-8)=0
W-10=0
W=10 ANS. FOR THE WIDTH.
L=18-10=8 ANS. FOR THE LENGTH.
W-8=0
W=8 ANS. FOR THE WIDTH.
L=18-8=10 ANS. FOR THE LENGTH.
PROOF:
10*8=80